OpenStudy (anonymous):

Given f(x) = –5x + 1, what is f(–2)? A. B. C. D.

OpenStudy (anonymous):

options?

OpenStudy (jadeishere):

Plug in 92 for x f(x) = -5(-2) + 1 f(x = 10 + 1 f(x) = 11

OpenStudy (jadeishere):

not 92, sorry -2

OpenStudy (anonymous):

a.1 b.-9 c.11 d.3/5

OpenStudy (anonymous):

when i divided 10 by 1 it equals 10???????????

OpenStudy (jadeishere):

?

OpenStudy (anonymous):

on a calculater 10/1=10

OpenStudy (jadeishere):

Um, yes? but where's the division sign?

OpenStudy (anonymous):

i dont have one on my keyboard

OpenStudy (jadeishere):

Oh so it's f(x) -5(-2) dvided by 1?

OpenStudy (anonymous):

yes

OpenStudy (jadeishere):

That doesn't make sense. I'm pretty sure it's 11 if there's no option that is 10

OpenStudy (anonymous):

ill try 11 thank you:]

OpenStudy (jadeishere):

yep :)